OpenStudy (anonymous):

Sum of n!/n^n, n=1 to infinity. Determine whether the series converges or diverges.

OpenStudy (zarkon):

if you start at n=1 (\(0^0\)) is not defined \[\frac{n!}{n^n}=\frac{(n)(n-1)\cdots(2)(1)}{(n)(n)\cdots(n)(n)}\] \[=\frac{(n)(n-1)\cdots(3)(2)}{(n)(n)\cdots(n)(n)}\cdot\frac{(2)(1)}{(n)(n)}\le\frac{2}{n^2}\]

OpenStudy (zarkon):

a simple comparison test gives the answer

OpenStudy (zarkon):

\[0\le \sum_{n=1}^{\infty}\frac{n!}{n^n}\le\sum_{n=1}^{\infty}\frac{2}{n^2}<\infty\]

OpenStudy (zarkon):

and again the sum should start at n=1 and not n=0

OpenStudy (anonymous):

Yeah I meant for it to say 1. Ill edit that now